# Trace anomaly contributions to semi-classical wormhole geometries

### v1

@@ IV Particle Trajectories Around the Wormhole
It is convenient to rewrite Eq. (71) in terms of the proper radial distance
$l(r)=\pm\int_{r_{0}}^{r}\frac{dr}{(1-b(r)/r)^{1/2}},$ (72)
which is finite for all finite values of $r$. Note that the spacetime is extended in such a way that $l$ monotonically increases from -$\infty$ to $\infty$. $l<0$ or $l>0$ correspond to two parallel universes joined together via a throat at $l=0$. Using the proper radial distance, Eq. (71) takes the simple form
@@
$l(r)=\pm r\sqrt{1-\frac{r_{0}}{r}}\pm\frac{r_{0}}{2}\ln\left[\frac{2r}{r_{0}}\left(\sqrt{1-\frac{r_{0}}{r}}+1\right)-1\right],$ (80)
@@
It can be readily shown that the second derivative of the effective potential at $\bar{r}_{2}$ is positive. In Fig. 6, we illustrate the behavior of the effective potential as a function of the proper radial distance for different values of the parameter $\alpha$. We observe that both the height and depth of the effective potential increase as $\alpha$ increases.
@@
$+$ $\left(\left(140\,{\xi}^{3}-280\,\xi\right){{\it r_{0}}}^{2}\left({r_{c}}-{r_{0}}\right)+\left(245\,{\xi}^{3}-35\,\xi\right){{r_{0}}}^{3}\right){{r_{c}}}^{9/2}+\left(-35\,{{r_{0}}}^{4}\xi+490\,{{r_{0}}}^{3}\xi\,\left({r_{c}}-{r_{0}}\right)\right){{r_{c}}}^{7/2}$
@@
$+\alpha\left[\left(280+420\,{\xi}^{2}\right){{r_{0}}}^{3}\sqrt{{r_{c}}-{r_{0}}}{{r_{c}}}^{4}+\left(-35\,{{r_{0}}}^{4}\xi+490\,{{\it r_{0}}}^{3}\xi\,\left({r_{c}}-{r_{0}}\right)\right){{r_{c}}}^{7/2}-210\,{{r_{c}}}^{3}{{\it r_{0}}}^{4}\sqrt{{r_{c}}-{r_{0}}}\right]=0.$
@@
$(\xi^{2}-1)r_{c}^{3}+(7-3\xi^{2})r_{0}r^{2}+\left(\frac{9}{4}\xi^{2}-15\right)r_{0}r+9r_{0}^{3}=0$ (90)

### v2

@@ IV Particle Trajectories Around the Wormhole
It is convenient to rewrite Eq. (77) in terms of the proper radial distance $l(r)$ which is finite for all finite values of $r$. Note that the spacetime is extended in such a way that $l$ monotonically increases from -$\infty$ to $\infty$ so that $l<0$ or $l>0$ correspond to two parallel universes joined together via a throat at $l=0$. Using the proper radial distance, Eq. (77) takes the simple form
@@
$l(r)=\pm r\sqrt{1-\frac{r_{0}}{r}}\mp\frac{r_{0}}{2}\ln\left[\frac{2r}{r_{0}}\left(\sqrt{1-\frac{r_{0}}{r}}+1\right)-1\right],$ (85)
@@
It can be readily shown that the second derivative of the effective potential at $\bar{r}_{2}$ is positive. In Fig. 6, we illustrate the behavior of the effective potential as a function of the proper radial distance for different values of the parameter $\alpha$. We observe that both the height and depth of the effective potential increase as $\alpha$ increases. At the throat a discontinuity in the derivative of the effective potential is observed which can be due to gluing two symmetric spacetimes through a thin shell, see e.g. [87] for more details. It is worth mentioning that for $\alpha=0$, the metric in Eq. (49) reduces to the wormhole solutions [84], whose gravitational lensing was investigated in Ref. [88].
@@
$+$ $\left(\left(140{\xi}^{3}-280\xi\right)r_{0}^{2}\left({r_{c}}-{r_{0}}\right)+\left(245{\xi}^{3}-35\,\xi\right)r_{0}^{3}\right)r_{c}^{9/2}+\left(-35r_{0}^{4}\xi+490r_{0}^{3}\xi\left({r_{c}}-{r_{0}}\right)\right)r_{c}^{7/2}$
@@
$+\alpha\left[\left(280+420{\xi}^{2}\right){{r_{0}}}^{3}\sqrt{{r_{c}}-{r_{0}}}{{r_{c}}}^{4}+\left(-35r_{0}^{4}\xi+490r_{0}^{3}\xi\left({r_{c}}-{r_{0}}\right)\right)r_{c}^{7/2}-210r_{c}^{3}r_{0}^{4}\sqrt{{r_{c}}-{r_{0}}}\right]=0.$
@@
$(\xi^{2}-1)r_{c}^{3}+(7-3\xi^{2})r_{0}r_{c}^{2}+\left(\frac{9}{4}\xi^{2}-15\right)r_{0}r_{c}+9r_{0}^{3}=0$ (95)